a coffee company wants a new flavor of Cajun coffee. How many pounds of coffee worth $10 a pound should be added to 30 pounds of coffee worth $3 a pound to get a mixture worth $5 a pound?
Answer by nerdybill(7384)

a coffee company wants a new flavor of Cajun coffee. How many pounds of coffee worth $10 a pound should be added to 30 pounds of coffee worth $3 a pound to get a mixture worth $5 a pound?
.
Let x = amount (lbs) of $10 per pound coffee
then
10x + 3(30) = 5(x+30)
10x + 30 = 5x+150
5x + 30 = 150
5x = 120
x = 24 pounds